Angle is 4 times measure of the complement. What's the measure of each angle? Please help and respond quickly:)
Answer by stanbon(75887)

Angle is 4 times measure of the complement. What's the measure of each angle?
------------
angle = x
complement = 90-x
----
Equation:
x = 4(90-x)
x = 360-4x
5x = 360
x = 72 degrees
90-x = 18 degrees
=======================
Cheers,
Stan H.